Find two consecutive even intergers such that the sum of their recipricols is 9/40.

Answer by checkley77(12844)

1/x+1/(x+2)=9/40
(x+2+x)/x(x+2)=9/40
(2x+2)/(x^2+2x)=9/40 cross multiply
9(x^2+2x)=40(2x+2)
9x^2+18x=80x+80
9x^2+18x-80x-80=0
9x^2-62x-80=0
(9x+10)(x-8)=0
9x-10=0
9x=-10
x=-10/9
x=-1.111 ans.
x-8=0
x=8 ans.